# AAAI Press Formatting Instructions for Authors Using — A Guide

### v2

@@ Introduction
Qwen2-vl Reader .83 .77 .68 .27 .60 .44 .48 .08 .34 .23 .59 .36
@@
Backbone Model VQA-RAD SLAKE PathVQA Mean
@@
Qwen2-vl Reader .73 .41 .84 .80 .87 .25 .81 .49
@@
Table 1: Results for classification (a) and visual question answering (b). Joint training consistently leads to the best results, with up to a 10-point increase in F1 score compared to fine-tuned (FT) RAG for multi-label classification. For visual question answering, we evaluate closed questions using the exact match metric and open questions using token recall. Using an oracle reranker demonstrates that even larger improvements are achievable with Jomed’s top-retrieved images in some datasets. Note that in the open question scenario, we do not evaluate the oracle because there is no clear discriminative boundary between different answers to determine the correct one, as there is in classification and closed question answering tasks. Using an oracle reranker shows that larger improvement is achievable with Jomed’s top retrieved images in some datasets.
Experiments
Medical
Pretrained
@@
Model Breast Derma
177K MedVInT-TE .88 .78
177K MedVInT-TD .90 .80
177K MedDrInternVL .72 -
177K MedDr + RADInternVL .88 -
255K GSCoInternVL .93 -
0 Jomed Qwen2-vl .87 .76
0 JomedPixtral .90 .80
Table 2: Accuracy of Jomed and prior reported results of medical pre-trained models for multiclass classification.
Medical
Pretrained
Size
Model VinDr-PCXR BRSET
255K MedDrInternVL .08 .08
255K GSCoInternVL .09 .33
0 Jomed Qwen2-vl .14 .37
0 JomedPixtral .14 .37
Table 3: F1 Comparison of Jomed to prior reported results of medical pre-trained models for multi-label classification.
Medical
Pretrained
Size
@@
Closed Open Closed Open Open
600K BiomedGPT-B .90 .85 .88 .28
600K LLaVA-MedLLaVA .83 .85 .91 .38
600K LLaVA-MedVicuna .85 .83 .92 .39
600K LLaVA-MedBioMedCLIP .87 .87 .91 .39
0 JomedQwen2-vl .90 .83 (.84) .93 .37 (.37)
0 JomedPixtral .90 .82 (.82) .90 .35 (.35)
@@
Table 4: Comparison of Jomed to medical pre-trained models for visual question answering. We report the token recall metric, reported for LLaVA-Med variants and token F1 reported for BiomedGPT (in red).
Experimental Setup
@@
Our datasets include real-world hospital datasets (BRSET (Nakayama et al. 2024) and VinDr‑PCXR (Pham, Tran, and Nguyen 2022)) alongside a variety of classification and visual question answering datasets. We focus on a low-resource data-efficient setting (training sets ranging from 546 to 7007 samples in classification, 1790-19,700 in VQA). Medical image annotation is a resource-intensive task, demanding expert annotators whose availability is limited and expensive. Clinical AI also often has poor generalization across heterogeneous medical centers and patient populations, often requiring that models be trained for the specific context in which they are deployed (Casey 2022). Such constraints often restrict researchers to datasets comprising only a few thousand samples for a given study. Retrieval augmentation is well-suited for this setting, as it is known to benefit low‑data regimes the most by leveraging external knowledge to compensate for sparse training samples.
@@
We consider three types of classification tasks: binary, multi-class, and multi-label. These include BreastMNIST (nickname: Breast; breast ultrasound imaging, binary) (Al-Dhabyani et al. 2020), DermaMNIST (nickname: Derma; pigmented skin lesion images, multi-class) (Tschandl, Rosendahl, and Kittler 2018), RetinaMNIST (nickname: Retina; retinal fundus images with multi-label disease annotations) (Liu et al. 2022), and two challenging real-world multi-label datasets: VinDr-PCXR (a chest X-ray dataset with 15 labels) (Pham, Tran, and Nguyen 2022) and BRSET (an ophthalmology dataset with 14 labels) (Nakayama et al. 2024). For visual question answering, we use widely adopted benchmarks: VQA‑RAD (Lau et al. 2018), SLAKE‑English (Liu et al. 2021), and PathVQA (He et al. 2020). Full details are provided in Appendix A.
@@
Retrieval augmentation is supported by an external index constructed from PubMed and medical records: PMC‑OA (Lin et al. 2023), MIMIC‑CXR (Johnson et al. 2019), and ROCO (Rückert et al. 2024). Full descriptions are available in Appendix A. Our experiments leverage two backbone models—Pixtral (12B) (Agrawal et al. 2024) and Qwen2‑vl (7B parameters) (Wang et al. 2024)—with jina‑clip‑v1 (Xiao, Mastrapas, and Wang 2024) serving as the general-purpose retriever’s visual head, embedding both the texts and images of the retrieval index. All models were trained in a low-resource setting, with all runs performed on a single L40s GPU. More details in Appendix B.
For comparison to medically pre‑trained LVLMs, we include competitive baseline models: BiomedGPT (Luo et al. 2023); three LLaVA-Med (Li et al. 2023) variants; two LVLM variants based on PMC CLIP (MedVInT-TE and MedVInT-TD) (Zhang et al. 2023b); and three variants based on InternVL (Chen et al. 2024): MedDr, MedDr + RAD (He et al. 2024a), and GSCo (He et al. 2024b). Full details are in Appendix C.
Results
Table 1(a) and Table 1(b) demonstrate the effectiveness of Jomed for two backbone models, Pixtral and Qwen2-vl, across five medical imaging classification benchmarks and three visual question answering benchmarks. Jomed achieved superior results compared to both the standard fine-tuned multimodal RAG baseline (trained only with retrieval augmentation fine-tuning, as detailed in Section Reader Retrieval Augmentation Fine-Tuning) and the Reader baseline (LVLM trained without retrieval), for both backbones and all benchmarks.
In Table 1(a), for classification tasks, Jomed with the Qwen2-vl backbone achieved a mean improvement of 0.05 in F1 and 0.03 in ACC over fine-tuned multimodal RAG, with F1 gains of 0.10 and 0.08 in two datasets. Jomed with Pixtral achieved a mean improvement of 0.03 in F1 and ACC. In Table 1(b), for VQA tasks, Jomed with Pixtral and Qwen2-vl achieved a mean improvement of 0.02 in ACC (Exact Match) for closed question answering. For open question answering, Jomed with Qwen2-vl achieved a mean improvement of 0.03 in token recall, while Jomed with Pixtral achieved an improvement of 0.02.
@@ Results
In addition, we evaluated our model in an oracle scenario. Instead of the final fusion of logits (Section Inference), we applied an oracle that, given the model responses for different retrieved images, selects the correct answer if it exists. Our findings in Tables 1(a) and 1(b) show that the correct answer exists in a large proportion of cases, yielding superior results. This shows that in these cases, Jomed is able to surface in its top retrieved images an image which could lead to a correct prediction, however the information gets lost in the process of prediction fusion with other retrieved images. Simply taking the label with highest confidence or label with highest average confidence across retrieved images, underperforms. This motivates us to explore the use of the o3 multimodal reasoning model to detect this image (see below).
Benchmarking versus pretrained LVLMs for classification. In Tables 2 and 3, we include previously reported results for several leading medically pretrained LVLM methods (detailed in Section Experimental Setup). Jomed with both backbones delivered competitive performance relative to models that underwent extensive pretraining. For multi-label classification, Jomed outperformed the pre-trained models. For binary and multi-class classification tasks, Jomed with Pixtral matched or exceeded all models except GSCo, which achieved 0.03 higher ACC on the Breast dataset. Overall, our models performed strongly across four medical benchmarks compared to medical pretrained models. We are not aware of any medical LVLM evaluated on Retina. To our knowledge, no other LVLMs have been evaluated on our benchmarks. Additionally, in Table 2 we report only ACC, as F1 scores were not reported for the compared models. Similarly, in Table 3, we report only F1 scores. We also evaluated a medical baseline, Med-Flamingo, with a pretrained medical retriever (BiomedCLIP (Johnson, Douze, and Jégou 2019)). We observed that Med-Flamingo generally underperformed, and showed limited improvement from utilizing the retrieved context to enhance performance, which we suspect may be attributed to its relatively older LLaMA backbone (Touvron et al. 2023). We further try BiomedCLIP as a retriever with our general-purpose LVLMs, which showed improvements over FT RAG, with a similar pattern to the general-purpose retriever. Full results in Appendix D.
Benchmarking versus pretrained LVLMs for visual question answering. In Table 4, we also compare our approach against medical pretrained LVLM methods (methods detailed in Section Experimental Setup) for VQA. Here too, Jomed delivered competitive performance relative to extensively pretrained models. For closed question tasks, Jomed with Pixtral and Qwen2-vl was superior to or matched current models. For open-ended question answering, Jomed with Qwen2-vl backbone matched LLaVA-MedVicuna on the SLAKE dataset and achieved slightly lower results for PathVQA, showing a 0.01 - 0.02 token recall gap compared to pretrained LVLMs. Note that we do not compare our method on VQA-RAD, since our approach is trained on an internal split of training, validation, and test sets, whereas VQA-RAD provides only official training and test splits.
@@
Table 5: Analysis of joint training effect: for inconsistent retrieval predictions, we can see that applying joint training boosts performance significantly for all datasets. For consistent retrieval predictions, the results are slightly improved. Note that inconsistent retrievals are challenging to all models, even non-retrieval ones.
@@
Table 6: Proportion of cases with inconsistent predictions, per dataset, for Jomed based on Qwen2-vl and Pixtral.
Analysis
@@
Both the consistent and inconsistent retrieval prediction sets are established after the initial reader training stage, and their performance is assessed following joint multimodal retriever fine-tuning. Detailed results are presented in Table 5. Overall, Jomed improves performance on inconsistent retrieval predictions by +0.12 in accuracy and +0.13 in F1 score when using the Qwen2‑vl backbone, and by +0.09 in both accuracy and F1 score with the Pixtral backbone, while also offering a slight improvement on the consistent.
We also observed that inconsistent retrieval predictions are empirically more challenging for all models (both with and without retrieval). There is approximately a 10–20 point performance gap between the inconsistent and consistent.
@@
Performance with o3 as reranker. We have observed that results for inconsistent predictions remain relatively low (Table 5), even after applying joint training (Table 5). However, in the oracle analysis reported above, we found that Jomed can sometimes retrieve images that lead to correct predictions, together with other images that lead to wrong predictions. We thus explore whether a state-of-the-art multimodal model, when used as an optional reranker, could close the gap toward oracle performance for inconsistent retrieval predictions. Specifically, we investigated whether the o3 reasoning model (OpenAI 2025) could effectively select the image from the retrieved set that offers the most predictive information, guiding the model toward correct classification. We provided o3 with an image to classify, along with four retrieved images+captions, and tasked it with identifying the one containing the most predictive information. (detailed prompt in Appendix D)
@@
Our results show that o3’s performance generally surpasses simply taking the image with highest confidence, but o3 is generally inferior compared to fusing logits in Jomed. Thus, we were unable to bridge the gap to oracle performance, and conclude that reranking in this setting remains a significant challenge even for a frontier model. We report full results in Appendix D.
Additional analyses and ablations. We include several robustness ablations and analyses in Appendix D. We observed that our model is robust to scenarios where the retrieval is noisy or not given to the models. Further, we observed that the retrieval itself includes predictive knowledge without being given the query. details in Appendix D.
Conclusion
We explored the effectiveness of a multimodal retriever jointly optimized with a general-purpose LVLM. Our lightweight approach yielded substantial improvements, including in challenging cases involving inconsistent retrieval predictions. Our oracle analysis revealed a considerable performance gap between actual results and what was theoretically achievable using the retrieved images, providing a foundation for future research on closing this gap.

### v3

@@ 3.3 Inference
LLaVA (7B) MMed-RAG .85 .84 .75 .30 .63 .46 .55 .11 .42 .30 .64 .40
Qwen2-vl (7B) Reader .83 .77 .68 .27 .60 .44 .48 .08 .34 .23 .59 .36
RAD .84 .79 .43 .34 .55 .40 .57 .09 .40 .25 .56 .37
@@
Backbone (Size) Model VQA-RAD SLAKE PathVQA Mean
@@
LLaVA (7B) MMed-RAG .74 .39 .87 .81 .90 .31 .84 .50
Qwen2-vl (7B) Reader .73 .41 .84 .80 .87 .25 .81 .49
@@
Table 1: Results for classification (a) and VQA (b) with non-medically pretrained LVLM backbones (see Tables 3, 4 for comparisons to with medical pre-training; MMed-RAG’s retrieval component is medically pre-trained). joint optimization consistently leads to the best results. For VQA, we evaluate closed questions using the exact match metric and open questions using token recall. Using an oracle reranker (for classification and closed VQA; §4.2) demonstrates that even larger improvements are achievable with Jomed’s top-retrieved images in some datasets.
MMed-RAG
LLaVaMed
variants
MedDr
variants
GSCo
MedVInT
variants
Medical Pre-Training
@@
600K 600K 255K 255K 177K
Table 2: Medical pre-training size for LVLMs under extensive medical pretraining.
4 Experiments
Model Breast Derma VinDr BRSET
MMed-RAGLLaVA-Med .89 .79 .11 .33
MedVInT-TE .88 .78 - -
MedVInT-TD .90 .80 - -
MedDrInternVL .72 - .08 .08
MedDr + RADInternVL .88 - - -
GSCoInternVL .93 - .09 .33
Jomed Qwen2-vl .87 .76 .14 .37
JomedPixtral .90 .80 .14 .37
Table 3: Comparison to prior reported results of medical pre-trained models, for binary and multiclass classification (Breast and Derma; accuracy), and for multi-label classification (VinDr-PCXR and BRSET; F1).
@@ 4 Experiments
Closed Open Closed Open
MMed-RAGLLaVA-Med .89 .84 .92 .39
BiomedGPT-B .90 .85 .88 .28
LLaVA-MedLLaVA .83 .85 .91 .38
LLaVA-MedVicuna .85 .83 .92 .39
LLaVA-MedBioMedCLIP .87 .87 .91 .39
JomedQwen2-vl .90 .84 (.84) .93 .38 (.37)
JomedPixtral .90 .84 (.82) .90 .36 (.35)
@@
Table 4: Comparison to medical pre-trained models for visual question answering. We report the token recall metric, reported for LLaVA-Med variants and token F1 reported for BiomedGPT (in red).
4.1 Experimental Setup
@@
Our datasets include real-world hospital datasets (BRSET Nakayama et al. (2024) and VinDr‑PCXR Pham et al. (2022)) alongside a variety of classification and visual question answering datasets. We focus on a low-resource data-efficient setting (training sets ranging from 546 to 7007 samples in classification, 1790-19,700 in VQA). Medical image annotation is a resource-intensive task, demanding expert annotators whose availability is limited and expensive. Clinical AI also often has poor generalization across heterogeneous medical centers and patient populations, often requiring that models be trained for the specific context in which they are deployed Casey (2022). Such constraints often restrict researchers to datasets comprising only a few thousand samples for a given study. Retrieval augmentation is well-suited for this setting, as it is known to benefit low‑data regimes the most by leveraging external knowledge.
@@
We consider three types of classification tasks: binary, multi-class, and multi-label. These include BreastMNIST (“Breast”; breast ultrasound imaging, binary) Al-Dhabyani et al. (2020), DermaMNIST (“Derma”; pigmented skin lesion images, multi-class) Tschandl et al. (2018), RetinaMNIST (“Retina”; retinal fundus images with multi-class disease annotations) Liu et al. (2022), and two challenging real-world multi-label datasets: VinDr-PCXR (a chest X-ray dataset with 15 labels) Pham et al. (2022) and BRSET (an ophthalmology dataset with 14 labels) Nakayama et al. (2024). For visual question answering, we use widely adopted benchmarks: VQA‑RAD Lau et al. (2018), SLAKE‑English Liu et al. (2021), and PathVQA He et al. (2020). Full details are provided in Appendix A.1.
@@
Retrieval augmentation is supported by an external index constructed from PubMed and medical records: PMC‑OA Lin et al. (2023), MIMIC‑CXR Johnson et al. (2019a), and ROCO Rückert et al. (2024). Full descriptions are available in Appendix A.2. Our experiments leverage two backbone models—Pixtral (12B) Agrawal et al. (2024) and Qwen2‑vl (7B parameters) Wang et al. (2024)—with jina‑clip‑v1 Xiao et al. (2024) serving as the general-purpose retriever’s visual head, embedding both the texts and images of the retrieval index. All runs were performed on a single L40s GPU. More details in Appendix B.
Baselines. We compare to several RAG baselines. RAD He et al. (2024a), a retrieval-based approach for classification tasks, where the label of the most similar training image is used as context. We integrate RAD with both backbone models (Qwen2-VL and Pixtral). MMed-RAG Xia et al. (2024a), a recent state-of-the-art multimodal RAG framework, using two backbones: LLaVA-Med and LLaVA, both paired with a medically pre-trained retriever (CLIP). We also adopt a standard fine-tuned RAG baseline in which the retriever is fixed and only the reader is fine-tuned, corresponding to the first phase of our approach. Finally, we compare our retriever loss with the base Perplexity Distillation Loss (PDist) Izacard et al. (2023), integrated into Jomed in place of our customized retrieval loss. Full implementation details are in Appendix C. For comparison to medically pre‑trained LVLMs, we include competitive baselines: BiomedGPT Luo et al. (2023); three LLaVA-Med Li et al. (2023) variants; two LVLM variants based on PMC CLIP (MedVInT-TE and MedVInT-TD) Zhang et al. (2023b); and three InternVLChen et al. (2024) variants: MedDr, MedDr + RAD He et al. (2024a), and GSCo He et al. (2024b). Details in Appendix C.
4.2 Results
Table 1(a) and Table 1(b) show the effectiveness of Jomed on five medical classification and three VQA benchmarks. We compare with methods of similar computational cost; all LVLMs were not extensively pre-trained on medical data, whereas the MMed-RAG retriever was. Jomed consistently outperforms MMed-RAG, RAD, the fine-tuned RAG baseline, Jomed + PDist loss, and the Reader baseline (LVLM without retrieval).
@@ 4.2 Results
In addition, we evaluated our model in an oracle scenario. Instead of the final fusion of logits (Section 3.3), we applied an oracle that, given the model responses for different retrieved images, selects the correct answer if it exists. [^2] For open-ended questions we do not conduct this analysis as there is no simple boundary between correct/incorrect. Our findings in Tables 1(a) and 1(b) show that the correct answer exists in a large proportion of cases, yielding superior results. This shows that in these cases, Jomed is able to surface in its top retrieved images an image which could lead to a correct prediction, however the information gets lost in the process of prediction fusion with other retrieved images. Simply taking the label with highest confidence or label with highest average confidence across retrieved images, underperforms. This motivates us to explore the use of the o3 multimodal reasoning model to detect this image. see Section 5.
@@
Table 5: Analysis of joint optimization effect: for inconsistent retrieval predictions, we can see that applying joint optimization boosts performance significantly for all datasets. For consistent retrieval predictions, the results are slightly improved. Note that inconsistent retrievals are challenging to all models, even non-retrieval ones.
@@
Table 6: Proportion inconsistent predictions cases, per dataset, for Jomed based on Qwen2-vl and Pixtral.
Benchmarking versus medically pretrained LVLMs for classification. In Table 3, we include previously reported results for several leading medically pretrained LVLM methods (detailed in Section 4.1). Jomed demonstrates competitive performance compared to models that underwent extensive medical pretraining (pretraining sizes summarized in Table 2). Jomed with the Pixtral backbone matches or exceeds all models except GSCo on the Breast dataset. To our knowledge, no medically pretrained LVLMs have been evaluated on the Retina benchmark, and we are not aware of any prior LVLM results on our other classification benchmarks. For binary and multi-class classification, we report only accuracy (ACC), as prior works did not report F1 scores. Conversely, for multi-label classification, we report only F1 scores. We also evaluated a medical baseline, Med-Flamingo, which generally underperformed and showed limited benefit from retrieval augmentation. We attribute this to its relatively older LLaMA backbone Touvron et al. (2023). Additionally, we experimented with using BiomedCLIP Johnson et al. (2019b) as the retriever for our general-purpose LVLMs, which exhibited a similar trend to the non-medical retriever. Full results in Appendix D.3.
Benchmarking against pretrained LVLMs for visual question answering. In Table 4, we compare our approach with medically pretrained LVLM methods (detailed in Section 4.1) for VQA. Jomed achieves competitive performance relative to extensively pretrained models. For closed-question tasks, Jomed outperforms or matches existing models. For open-ended question answering, Jomed with the Qwen2-vl backbone matches LLaVA-MedVicuna on the SLAKE dataset and LLaVA-MedLLaVA on the PathVQA dataset. Note that we do not report results on VQA-RAD, since our method was trained using an internal split of training and validation, whereas VQA-RAD provides only official training and test splits.
5 Analysis
@@
Both the consistent and inconsistent retrieval prediction sets are established after the initial reader training stage, and their performance is assessed following joint multimodal retriever fine-tuning. Detailed results are presented in Table 5. Overall, Jomed improves performance on inconsistent retrieval predictions by +0.12 in accuracy and +0.13 in F1 score when using the Qwen2‑vl backbone, and by +0.09 in both accuracy and F1 score with the Pixtral backbone, while also offering a slight improvement on the consistent. We also observed that inconsistent retrieval predictions are empirically more challenging for the reader, FT RAG and Jomed models, with a 10–20 point performance gap between the inconsistent and consistent.
@@
Performance with o3 as reranker. Results for inconsistent predictions remain relatively low (Table 5), even after applying joint optimization (Table 5). However, in the oracle analysis reported above, we found that Jomed can sometimes retrieve images that lead to correct predictions, together with other images that lead to wrong predictions. We thus explore whether a state-of-the-art multimodal model, when used as an optional reranker, could close the gap toward oracle performance. Specifically, we investigated whether the o3 reasoning model OpenAI (2025) could effectively select the image from the retrieved set that offers the most predictive information, guiding the model toward correct classification. We provided o3 with an image to classify, along with four retrieved images+captions, and tasked it with identifying the one containing the most predictive information.
@@
Our results show that o3’s performance generally surpasses simply taking the image with highest confidence, but o3 is generally inferior compared to fusing logits in Jomed.We conclude that reranking in this setting remains a significant challenge even for a frontier model. Results in Appendix D.2.
Additional analyses and ablations. We evaluated the contribution of each component in our method in Appendix D.1. Additionally, we present several robustness ablations in Appendix D.4. We demonstrate robustness across different scenarios.
6 Conclusion
We demonstrated that joint optimization of general-purpose LVLMs and multimodal retrievers can achieve competitive medical diagnosis performance through lightweight fine-tuning, without medical pre-training. Future work may explore whether this approach generalizes as a cost-effective alternative to domain-specific pre-training when task-specific data and domain knowledge bases are available but large-scale pre-training is prohibitive. In addition, we identified and substantially mitigated inconsistent retrieval predictions, where different retrieved candidates lead to conflicting diagnoses. Our oracle analysis revealed a considerable performance gap between actual results and what was theoretically achievable using the retrieved images, providing a foundation for future research on closing this gap. Future work may also look into the prevalence and impact of inconsistent retrieval predictions more broadly.
Limitations
Sequential rather than end-to-end optimization.
We acknowledge that, in our method, gradients do not flow simultaneously between the retriever and reader modules. Nevertheless, the reader leverages signals from the retriever during training to improve its performance, while the retriever, in turn, benefits from feedback provided by the reader to refine its retrieved content. This bidirectional interaction effectively induces a joint optimization effect within our framework.
Scope of evaluation.
Our evaluation focuses on classification and visual question-answering tasks. We did not evaluate our method on report generation, which requires different capabilities (longer-form generation, clinical writing conventions) and would benefit from dedicated investigation. Additionally, while our method shows consistent improvements across diverse imaging modalities (ultrasound, fundus photography, X-ray, histopathology, dermatoscopy, CT) and tasks (visual question answering, malignant lesion classification, diabetic retinopathy severity grading), evaluation on additional specialized modalities (e.g., PET, Microscopy, OCT) and tasks (e.g., organ classification, retinal OCT disease classification) would further validate generalizability.
@@ Appendix A Dataset
A.1 Dataset for Evaluation
Our datasets include real-world hospital datasets (BRSET Nakayama et al. (2024) and VinDr‑PCXR Pham et al. (2022)) alongside a variety of classification and visual question answering datasets. We focus on a low-resource data-efficient setting (training sets ranging from 546 to 7007 samples in classification, 1790-19,700 in VQA). Medical image annotation is a resource-intensive task, demanding expert annotators whose availability is limited and expensive. Clinical AI also often has poor generalization across heterogeneous medical centers and patient populations, often requiring that models be trained for the specific context in which they are deployed Casey (2022). Such constraints often restrict researchers to datasets comprising only a few thousand samples for a given study. Retrieval augmentation is well-suited for this setting, as it is known to benefit low‑data regimes the most by leveraging external knowledge to compensate for sparse training samples. For the classification tasks, we used the following datasets:
BreastMNIST (nickname: Breast) Al-Dhabyani et al. (2020) is a binary classification dataset of breast ultrasound. Following the official split, we use 546 for training, 78 for validation and 156 for testing.
RetinaMNIST (nickname: Retina) Liu et al. (2022) is a multi-label classification dataset of retina Fundus Camera. Following the official split, we use 1,080 for training, 120 for validation and 400 for testing.
DermaMNIST (nickname: Derma) Tschandl et al. (2018) is a multi-class classification dataset of common pigmented skin lesions. Following the official split, we use 7,007 for training, 1,003 for validation and 2,005 for testing. The dataset contains clinical images from 7 different diagnostic categories: actinic keratoses, basal cell carcinoma, benign keratosis, dermatofibroma, melanoma, melanocytic nevi, and vascular lesions.
VinDr-PCXR Pham et al. (2022) is a large, open pediatric chest X-ray dataset collected in Vietnam (2020–2021) containing 9,125 posteroanterior scans from patients under 10 years old. It provides both lesion-level bounding-box annotations for 36 findings and image-level labels for multi-label of 15 diagnoses, curated by experienced radiologists. We follow the official split, which includes 7,728 training and 1,397 test images, with an additional internal split of the training data into 85% for training and 15% for validation.
BRSET Nakayama et al. (2024) is a Brazilian multilabel ophthalmological dataset comprising retinal fundus images annotated with 14 distinct pathological findings. The dataset, composed of 16,266 images, presents a challenging real-world scenario for multilabel classification. There is no publicly available split; therefore, we apply an internal split for train, validation, and test sets (70% train, 10% validation, and 20% test).
Breast, Derma, and Retina taken from the large-scale MNIST-like collection of standardized biomedical images, including 12 datasets for 2D and 6 datasets for 3D. We especially used MedMNIST+ Yang et al. (2023), which is a higher resolution extension of the original MedMNIST. We used the highest resolution of 224 × 224.
For medical question answering, we have three well-known datasets:
VQA-RAD Lau et al. (2018) is a clinician-annotated visual question answering dataset focused on radiology images (e.g., X-ray, CT, MRI). It pairs each image with both open-ended and yes/no questions. Following the official split, we use 1,753 questions for training and 453 for testing. In addition, we further partition the training set into 85% for training and 15% for validation.
PathVQA He et al. (2020) is a pathology-focused medical VQA dataset built from textbooks and the PEIR digital library, comprising 4,289 images and 32,632 question–answer pairs spanning both open-ended and yes/no questions. We follow the official split provided by the authors: 19,700 for training, 6,260 for validation, and 6,720 for testing.
SLAKE Liu et al. (2021) is a bilingual medical VQA dataset. We use the English subset, SLAKE-English, which comprises 642 radiology images and 7.03k English QA pairs. We follow the official train/validation/test split of 4.92k/1.05k/1.06k QA pairs, and include all question types, covering both open-ended and closed-ended formats.
A.2 Dataset for Index
For construction of the Index we used three large datasets of (image, text) pairs: PMC-OA Lin et al. (2023), ROCO Rückert et al. (2024) and MIMIC-CXR Johnson et al. (2019a):